# Demystifying Multilingual Chain-of-Thought in Process Reward Modeling

### v1

@@ 1 Introduction
Aligning large language models (LLMs) with human preferences can significantly improve the model performance across various downstream tasks (Christiano et al., 2017; Ziegler et al., 2019). This requires a reward model that is trained on human preference data (Ziegler et al., 2019; Stiennon et al., 2020; Shen et al., 2021; Ouyang et al., 2022). Typically, reward models are trained based on the final outcome of the LLMs’ response, and we refer to these as outcome reward models (ORMs) (Cobbe et al., 2021a; Uesato et al., 2022; Yu et al., 2023a). However, most of recent work demonstrates that ORMs fall short on complex multi-step reasoning tasks (Uesato et al., 2022; Shao et al., 2024). To overcome this limitation, process reward models (PRMs) are introduced, providing fine-grained rewards at each step of the LLMs’ chain-of-thought (Lightman et al., 2024; Li et al., 2023; Wang et al., 2024b; Ma et al., 2023). Previous research has shown that LLMs supervised by PRMs can effectively produce better responses (Wang et al., 2024b; Shao et al., 2024).
@@
In this work, our main takeaways can be summarized as follows:
@@
  • Multilingual PRM can benefit even more from more candidate responses and trainable parameters. Our analysis demonstrates that PRM-multi becomes even more advantageous when there is a larger number of candidate responses (see Section 6.2) and when more trainable parameters are introduced (see Section 6.3).
@@ 3.1 PRM Training
Given a question $p$ and its solution $s$, the ORM assigns a single value to $s$ to indicate whether $s$ is correct. We stack a binary classifier on top of the LLM and train the ORM with the binary cross-entropy loss:
@@ 4 Experimental Setups
We combine the PRM800K (Lightman et al., 2024) and Math-Shepherd (Wang et al., 2024b) as training data to finetune PRMs, and translate the combined dataset from English (en) to six languages: German (de), Spanish (es), French (fr), Russian (ru), Swahili (sw), and Chinese (zh) with using NLLB 3.3B (Costa-jussà et al., 2022). The reasoning step statistics are presented in Table 4 (Appendix A), and the parallel examples across seven languages have the same number of reasoning steps.
@@
We evaluate the performance of LLMs using two widely used math reasoning datasets, MGSM (Shi et al., 2022) and MATH500 (Wang et al., 2024b). For the MATH500 datset, we translate it from English to ten languages: Bengali (bn), German (de), Spanish (es), French (fr), Japanese (ja), Russian (ru), Swahili (sw), Telugu (te), Thai (th), and Chinese (zh) with Google Translate, which is consistent with the languages included in the MGSM dataset. Furthermore, we also categorize the languages involved in the downstream tasks into two groups based on the training data of PRM: seen languages (en, de, es, fr, ru, sw, and zh) and unseen languages (bn, ja, te, and th).
Multilingual PRM Setups
To better understand PRMs in the context of multilingual research, we define three setups: PRM-mono, PRM-cross, and PRM-multi. The PRM-mono setup is trained and evaluated on the same single language, serving as the baseline for monolingual PRMs. The PRM-cross setup is trained on one language but evaluated on all 11 test languages. Specifically, in this work, we train PRM-cross on the English PRM dataset unless otherwise specified. Finally, the PRM-multi setup represents the multilingual PRM, which is both trained on all the seen languages and evaluated on all 11 test languages. To enhance the reliability and generalizability of our study, we train our multilingual PRM (verifier) based on the Qwen2.5-Math-7B-Instruct (Yang et al., 2024a), and leverage three diverse LLMs as the generator: MetaMath-Mistral-7B (Yu et al., 2023b), Llama-3.1-8B-math (fine-tuned with the MetaMath dataset (Dubey et al., 2024)), [^2] https://huggingface.co/gohsyi/Meta-Llama-3.1-8B-sft-metamath and DeepSeekMath-7B-Instruct (Shao et al., 2024). The details of training these PRMs are presented in Appendix B.
@@ 5.1 Monolingual, Cross-lingual, or Multilingual PRMs?
Building upon the findings of Wu et al. (2024b), who demonstrated that cross-lingual ORMs outperform monolingual ones, we investigate the impact of multilingualism on PRMs. Specifically, we compare PRM-mono, PRM-cross, and PRM-multi to determine which setup offers best performance across languages.
@@
As shown in Table 1, PRM-multi consistently achieves the highest performance across multiple language generators on the seen languages, surpassing PRM-mono and PRM-cross by +1.5 and +1.2 with Llama-3.1-8B-math generator, respectively. This indicates that incorporating data from multiple languages for PRM training significantly enhances the model’s ability across different languages. When comparing PRM-mono and PRM-cross, we observe that PRM-cross outperforms the PRM-mono for the English-centric MetaMath-Mistral-7B and Llama-3.1-8B-math generators. We hypothesize that this advantage stems from the pre-training phase: these generators are predominantly trained on English data but have limited exposure to multilingual corpora. As a result, fine-tuning on English PRM data enhances the reasoning capabilities of PRMs, facilitating greater cross-lingual transfer. More monolingual results are in Appendix C.
@@
Both PRM-cross and PRM-multi are evaluated on four additional unseen languages. As shown in Table 1, PRM-multi demonstrates superior overall performance on the unseen languages in terms of $\mu_{\textsc{unseen}}$. These results suggest that training PRMs on multilingual datasets can effectively enhance model generalization to the unseen languages.
In conclusion, these findings demonstrate that training a single multilingual PRM is an effective strategy for broad cross-lingual coverage, outperforming models trained either on a target language or on English alone. This outcome supports that PRM-multi is particularly advantageous for expanding the capabilities of PRMs in multilingual settings. More results on MGSM are in Appendix D.
@@ 5.2 Does More Languages Lead to Better Multilingual PRMs?
We conduct experiments by training PRMs on datasets ranging from a single language up to all seven languages. In this section, the number of total training examples of all PRMs are fixed. When the number of languages exceeds one, the total training examples are evenly distributed across all the selected languages. For evaluation, we test all PRMs on 11 different languages. The evaluation scores are averaged for each test language across all PRMs trained with the same number of languages.
@@ 5.3 How Much English Data Do We Need for Multilingual PRMs?
Figure 3: Best-of-N sampling performance of Llama-3.1-8B-math with PRMs finetuned on a training set where Pof the data is in English and (100 - P)is uniformly distributed across six other languages. Each tick on the X-axis represents a specific tuning set configuration. The dash lines in blue, red, and green, indicate the average scores of all the languages, the seen languages, and the unseen languages, respectively.
@@ 6.1 Which Steps Are More Prone to Errors?
PRMs provide fine-grained feedback on each intermediate step of a model’s chain-of-thought reasoning process. The errors made at intermediate steps can propagate through the reasoning chain, ultimately affecting the final answer. Therefore, in this section, we investigate the earliest errors made by PRMs during the reasoning process, following Zheng et al. (2024).
@@ 6.2 Do More Candidates Drive Better Performance?
Recent research suggests that providing more candidate solutions can significantly boost the performance of PRM (Lightman et al., 2024; Wang et al., 2024b, a). To investigate whether this trend extends to multilingual settings, we examine the impact of varying the number of candidates on PRM-mono, PRM-cross, and PRM-multi.
@@
We conduct experiments on the MATH500 benchmark using the Llama-3.1-8B-math generator to compare the performance of multilingual PRM (PRM-multi), cross-lingual PRM (PRM-cross), and monolingual PRM (PRM-mono). For each approach, we vary the number of candidates N from 2 to 64. This allows us to assess how the number of candidate solutions influences performance across different PRM strategies in a multilingual context.
@@ 6.4 Does PRM Surpass ORM in the Multilingual Scenario?
In this section, we explore whether PRM also outperforms Outcome Reward Model (ORM) and self-consistency (SC) (Wang et al., 2022) in multilingual settings.
@@
Following Lightman et al. (2024); Wang et al. (2024b), we evaluate the performance of PRM-multi by comparing it with other verifier methods, including: Direct prediction (Baseline), Self-consistency (majority voting) (SC), and ORM. The accuracy of the best-of-N solution is used as the evaluation metric. Specifically, we train a multilingual ORM using uniform example budgets across seven seen languages. Then we assess the performance of verifiers on seven seen languages as well as on four additional unseen languages.
@@ 6.5 Can Multilingual PRM Enhance LLMs?
We have previously shown that multilingual PRM can bolster model performance under a best-of-N selection framework. In this section, we demonstrate that the multilingual PRM can be used as the reward model for finetuning the LLMs under a reinforcement learning paradigm.
@@
The results shown in Table 3 indicate that step-by-step PPO with PRM-multi (PPO-PRM) consistently outperforms a standard supervised fine-tuned Baseline and PPO with ORM (PPO-ORM). Llama-3.1-8B-math with PPO-PRM achieves average boosts of +1.86 and +2.19 across 11 languages, compared to Baseline and PPO-ORM, respectively. These findings highlight the importance of fine-grained multilingual reward signals. These gains demonstrate that process rewards can refine policy decisions for both reasoning steps and final outputs with reinforcement learning.
@@ 7 Conclusion
Our work demonstrates that multilingual PRMs significantly enhance the ability to perform complex, multi-step reasoning tasks in various languages, consistently outperforming both monolingual and cross-lingual counterparts. This conclusion is supported by comprehensive evaluations spanning 11 languages. Furthermore, our findings highlight that performance is sensitive to the number of languages and the volume of English training data. However, it also benefits substantially from more candidate responses and model parameters. These results underscore the importance of diverse language training in providing fine-grained rewards and open up promising avenues for multilingual reasoning.
@@ Appendix A Data Statistics
The dataset statistics are summarized in Table 4. These include the total number of examples, as well as the maximum, minimum, and average number of reasoning steps in the answers across all examples.
Appendix B Training Details
@@
Appendix C Cross-lingual Transfer of PRMs
@@
Appendix D Breakdown Results of MGSM for PRM-mono, PRM-cross, and PRM-multi

### v2

@@ 1 Introduction
Aligning large language models (LLMs) with human preferences can significantly improve the model performance across various downstream tasks (Christiano et al., 2017; Ziegler et al., 2019). This requires a reward model that is trained on human preference data (Ziegler et al., 2019; Stiennon et al., 2020; Shen et al., 2021; Ouyang et al., 2022). Typically, reward models are trained based on the final outcome of the LLMs’ response, and we refer to these as outcome reward models (ORMs) (Cobbe et al., 2021a; Uesato et al., 2022; Yu et al., 2023a). However, most of recent work demonstrates that ORMs fall short on complex multi-step reasoning tasks (Uesato et al., 2022; Shao et al., 2024). To overcome this limitation, process reward models (PRMs) are introduced, providing fine-grained rewards at each step of the LLMs’ reasoning (Lightman et al., 2024; Li et al., 2023; Wang et al., 2024b; Ma et al., 2023). Previous research has shown that LLMs supervised by PRMs can effectively produce better responses (Wang et al., 2024b; Shao et al., 2024).
@@
Our main takeaways are summarized as follows:
@@
  • Multilingual PRM can benefit more from more candidate responses and trainable parameters. Our analysis demonstrates that PRM-multi becomes more advantageous with a larger number of candidate responses (see Section 6.2) and when more trainable parameters are introduced (see Section 6.3).
@@ 3.1 PRM Training
Given a question $p$ and its solution $s$, the ORM assigns a single value to $s$ to indicate if $s$ is correct. We stack a binary classifier on top of the LLM and train the ORM with the binary cross-entropy loss:
@@ 4 Experimental Setups
We combine the PRM800K (Lightman et al., 2024) and Math-Shepherd (Wang et al., 2024b) as training data to finetune PRMs, and translate the combined dataset from English (en) to six languages: German (de), Spanish (es), French (fr), Russian (ru), Swahili (sw), and Chinese (zh) with using NLLB 3.3B (Costa-jussà et al., 2022). The reasoning step statistics are presented in Table 4 (Appendix A), and the parallel examples across seven languages have the same number of reasoning steps. To ensure high translation quality, we use regular expressions to filter out translated training instances that contain discrepancies in numbers or equations compared to the original English dataset.
@@
We evaluate the performance of LLMs using two widely used math reasoning datasets, MGSM (Shi et al., 2022) and MATH500 (Wang et al., 2024b). For the MATH500 datset, we translate it from English to ten languages: Bengali (bn), German (de), Spanish (es), French (fr), Japanese (ja), Russian (ru), Swahili (sw), Telugu (te), Thai (th), and Chinese (zh) with Google Translate, which is consistent with the languages included in the MGSM dataset. Furthermore, we also categorize the languages involved in the downstream tasks into two groups based on the training data of PRM: seen languages (en, de, es, fr, ru, sw, and zh) and unseen languages (bn, ja, te, and th). To ensure the quality of our testset, we employ two human translators to post-edit the translated examples for each high-resource language (de, es, fr, ru, zh, and ja) and leverage gpt-4o to revise the translations in low-resource languages (bn, sw, te, and th). More details are shown in Appendix B.
@@
Multilingual PRM Setups
To better understand PRMs in the context of multilingual research, we define three setups: PRM-mono, PRM-cross, and PRM-multi. The PRM-mono setup is trained and evaluated on the same single language, serving as the baseline for monolingual PRMs. The PRM-cross setup is trained on one language but evaluated on all 11 test languages. Specifically, in this work, we train PRM-cross on the English PRM dataset unless otherwise specified. Finally, the PRM-multi setup represents the multilingual PRM, which is both trained on all the seen languages and evaluated on all 11 test languages. To enhance the reliability and generalizability of our study, we train our multilingual PRM (verifier) based on the Qwen2.5-Math-7B-Instruct (Yang et al., 2024a), and leverage three diverse LLMs as the generator: MetaMath-Mistral-7B (Yu et al., 2023b), Llama-3.1-8B-math (fine-tuned with the MetaMath dataset (Dubey et al., 2024)), [^2] https://huggingface.co/gohsyi/Meta-Llama-3.1-8B-sft-metamath and DeepSeekMath-7B-Instruct (Shao et al., 2024). The details of training these PRMs are presented in Appendix C.
@@ 5.1 Monolingual, Cross-lingual, or Multilingual PRMs?
Building on Wu et al. (2024b)’s findings that cross-lingual ORMs outperform monolingual ones, we investigate the impact of multilingualism on PRMs. Specifically, we compare PRM-mono, PRM-cross, and PRM-multi to determine which setup offers best performance across languages.
@@
As shown in Table 1, PRM-multi consistently achieves the highest performance across multiple language generators on the seen languages, surpassing PRM-mono and PRM-cross by +1.5 and +1.2 with Llama-3.1-8B-math generator, respectively. This indicates that incorporating data from multiple languages for PRM training significantly enhances the model’s ability across different languages. When comparing PRM-mono and PRM-cross, we observe that PRM-cross outperforms the PRM-mono for the English-centric MetaMath-Mistral-7B and Llama-3.1-8B-math generators. We hypothesize that this advantage stems from the pre-training phase: these generators are predominantly trained on English data but have limited exposure to multilingual corpora. As a result, fine-tuning on English PRM data enhances the reasoning capabilities of PRMs, facilitating greater cross-lingual transfer. More monolingual results are in Appendix D.
@@
Both PRM-cross and PRM-multi are evaluated on four additional unseen languages. As shown in Table 1, PRM-multi demonstrates superior overall performance on the unseen languages in terms of $\mu_{\textsc{unseen}}$. These results suggest that training PRMs on multilingual datasets can effectively enhance model generalization to the unseen languages. More results on general-purpose LLM are provided in Appendix F.
In conclusion, these findings demonstrate that training a single multilingual PRM is an effective strategy for broad cross-lingual coverage, outperforming models trained either on a target language or on English alone. This outcome supports that PRM-multi is particularly advantageous for expanding the capabilities of PRMs in multilingual settings. More results on MGSM are in Appendix E.
@@ 5.2 Does More Languages Lead to Better Multilingual PRMs?
We conduct experiments by training PRMs on datasets ranging from a single language up to all seven languages.
In this section, the number of total training examples of all PRMs are fixed. When the number of languages exceeds one, the total training examples are evenly distributed across all the selected languages. For evaluation, we test all PRMs on 11 different languages. The evaluation scores are averaged for each test language across all PRMs trained with the same number of languages.
@@ 5.3 How Much English Data Do We Need for Multilingual PRMs?
Figure 3: Best-of-N sampling performance of Llama-3.1-8B-math with PRMs finetuned on a training set where P% of the data is in English and (100 - P)% is uniformly distributed across six other languages. Each tick on the X-axis represents a specific tuning set configuration. The dash lines in blue, red, and green, indicate the average scores of all the languages, the seen languages, and the unseen languages, respectively.
@@ 6.1 Which Steps Are More Prone to Errors?
PRMs provide fine-grained feedback on each intermediate step of a model’s chain-of-thought reasoning process. Errors at intermediate steps can propagate through the reasoning chain, ultimately affecting the final answer. Therefore, we investigate the earliest errors made by PRMs during the reasoning process, following Zheng et al. (2024).
@@ 6.2 Do More Candidates Drive Better Performance?
Recent research suggests that providing more candidate solutions can significantly boost the performance of PRM (Wang et al., 2024a, b). To explore if this applies in multilingual settings, we examine the impact of varying the number of candidates on PRM-mono, PRM-cross, and PRM-multi.
@@
We conduct experiments on the MATH500 benchmark using the Llama-3.1-8B-math generator to compare the performance of PRM-multi, PRM-cross, and PRM-mono. For each approach, we vary the number of candidates N from 2 to 64. This allows us to assess how the number of candidate solutions influences performance across different PRM strategies in a multilingual context.
@@ 6.4 Does PRM Surpass ORM in the Multilingual Scenario?
In this section, we explore whether PRM also outperforms Outcome Reward Model (ORM) and self-consistency (SC) in multilingual settings.
@@
Following Lightman et al. (2024); Wang et al. (2024b), we evaluate the performance of PRM-multi by comparing it with other verifier methods, including: Direct prediction (Baseline), Self-consistency (majority voting) (SC), and ORM. Specifically, we train a multilingual ORM using uniform example budgets across seven seen languages. Then we assess the performance of verifiers on seven seen languages as well as on four additional unseen languages.
@@ 6.5 Can Multilingual PRM Enhance LLMs?
In this section, we demonstrate that the multilingual PRM can be used as the reward model for finetuning the LLMs under a RL paradigm.
@@
The results shown in Table 3 indicate that step-by-step PPO with PRM-multi (PPO-PRM) consistently outperforms a standard supervised fine-tuned Baseline and PPO with ORM (PPO-ORM). Llama-3.1-8B-math with PPO-PRM achieves average boosts of +1.86 and +2.19 across 11 languages, compared to Baseline and PPO-ORM, respectively. These findings highlight the importance of fine-grained multilingual rewards. These gains demonstrate that process rewards can refine policy decisions for both reasoning steps and final outputs with RL. More results are in Appendix G.
@@ 7 Conclusion
Through comprehensive evaluations spanning 11 languages, our work demonstrates that multilingual PRMs significantly enhance the ability to perform complex, multi-step reasoning tasks in various languages, consistently outperforming both monolingual and cross-lingual counterparts. Furthermore, our findings highlight that PRM performance is sensitive to the number of languages and the volume of English training data. The multilingal PRMs also benefit from more candidate responses and model parameters. These results underscore the importance of diverse language training in providing fine-grained rewards and open up promising avenues for multilingual reasoning.
@@ 8 Limitations
Acknowledgement
This work is funded by EU Horizon Europe (HE) Research and Innovation programme grant No 101070631, and UK Research and Innovation under the UK HE funding grant No 10039436.
The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
@@ Appendix A Data Statistics
The dataset statistics are summarized in Table 4. These include the total number of examples, as well as the maximum, minimum, and average number of reasoning steps in the answers across all examples. For the selection criteria for the six languages, there are two key desiderata for the language selection in our work. Firstly, the examples must be accurately translatable into the target language by MT systems. Secondly, the target language must allow for proper evaluation. With these desiderata in mind, we selected six high-resource languages covered by the MGSM dataset. This choice ensures that the translated data closely aligns with the original English dataset and allows us to focus on comparing model strategies without introducing the added variability that lower-resource language translations might cause. We will clarify this in our future revision.
Appendix B Translation Details
Due to imbalanced resources across languages, translation has become a standard method for multilingual research. Recent research has demonstrated that machine-translated datasets are comparable to human-translated ones and can be directly used for training and evaluation (Chen et al., 2024; Thellmann et al., 2024).
In this study, after translating the English dataset into foreign languages, we use regular expressions to filter out the translated training instances that contain discrepancies in numbers or equations compared to the original English dataset. This ensures the correctness of the mathematical content. For the translated multilingual MATH500 test set, we employ two human translators to post-edit the test instances in high-resource languages (de, es, fr, ru, zh, and ja) by correcting inaccurate translations and verifying the consistency of mathematical notations. We pay \$0.05 USD for each example, resulting in a total cost of \$150 USD for post-editing. For the low-resource languages (bn, sw, te, and th) in MATH500, we leverage gpt-4o to post-edit the translations.
To verify the quality of our translations, we use Google Translate to back-translate the multilingual MATH500 and 1,000 random training instances from each training set into English. We then calculate the BLEU score using the original English instances as the reference translation. As shown in Table 5, the high BLEU scores confirm the quality of the translations in our datasets.
de es fr ru sw zh ja bn te th
Train 81.2 88.4 87.3 74.0 87.3 80.8 - - - -
Test 85.9 91.5 91.0 73.0 90.3 84.4 84.3 65.3 65.8 80.9
Table 5: BLEU scores of back-translation examples with using the original English data.
Appendix C Training Details
@@
Appendix D Cross-lingual Transfer of PRMs
@@
Appendix E Breakdown Results of MGSM for PRM-mono, PRM-cross, and PRM-multi
@@
Appendix F Results on General-Purpose LLM
We provide the results of the general LLM Qwen2.5-7B-Instruct on MATH500 in Table 7. It can be observed that the multilingual PRM achieves consistent conclusions when applied to the general LLM.
Baseline 36.3 37.9 33.4 44.2 40.6 41.6 40.2 40.8 20.8 37.4 40.6 38.2 20.2 34.4
PRM-mono - 54.2 - 61.0 57.6 58.0 57.0 58.2 31.6 56.2 - - -
PRM-cross 53.2 57.1 55.7 63.6 60.8 60.4 60.4 61.6 33.8 59.4 60.8 58.8 36.8 56.4
PRM-multi 54.0 58.2 56.7 64.8 61.6 61.8 61.2 62.2 35.2 60.6 61.2 58.6 38.2 57.8
Table 7: Performance on general LLM Qwen2.5-7B-Instruct.
Checkpoint Bengali English French
Baseline 45.2 78.4 67.6
Checkpoint-500 46.8 80.0 69.2
Checkpoint-1150 (final) 40.4 82.4 71.6
Table 8: The influence of final checkpoint selection strategy during the PPO training process.
Appendix G Influence of Checkpoint Selection
We observe a decline in Bengali performance in both ORM and PRM, as shown in Table 3. Upon evaluating the performance at each intermediate checkpoint, our analysis indicates that this behavior stems from the PPO training process and the strategy used for selecting the final checkpoint, as illustrated in Table 8. Specifically, since the checkpoint is selected based on the average loss across all languages, the one that minimizes the overall loss does not necessarily yield optimal performance for individual languages. In this case, Bengali appears to follow a distinct learning rate trajectory compared to other languages. We acknowledge this limitation and plan to investigate language-specific adjustments to the training process in future work.